OpenStudy (anonymous):

2x^2+4x-3 in the form of f(x)=a(x-h)^2+k

OpenStudy (nowhereman):

Those two forms are equal if the coefficients are the same. So look at \[2x^2 + 4x - 3 = ax^2 - 2ahx-h^2+k\] identify the coefficients and solve for a, h and k.

OpenStudy (radar):

I am treading on dangerous ground here. When i substitute the coefficients I come up with a=2, h=1 and k=-3\[f(x)=2(x-1)^{2}-3\] Not 100% sure about it though. What have you come up with?

OpenStudy (radar):

the k is wrong it has to be -5

OpenStudy (radar):

Hopefully nowhereman will bail me out of this quagmire!

OpenStudy (nowhereman):

To finish this, you get (sorry should be +a in front of h^2 up there): \[2 = a,\quad 4 = -2ah,\quad -3 = ah^2 + k\] which lead to \[ a = 2,\quad h = -1,\quad k = -5\] so that you have \[2x^2 + 4x - 3 = 2(x+1)^2-5\]